# reference.bib

### v1

@@ Abstract.
[…]
@@ Acknowledgements.
The author was supported by the HIMR/UKRI "Additional Funding Programme for Mathematical Sciences", grant number EP/V521917/1 as well as by a scholarship by Onassis Foundation. The author would also like to thank Prof. Thanasis Bouganis for suggesting this problem as well as for his continuous guidance and support.
@@ 1. Introduction
Kohnen and Skoruppa in [kohnen_skoruppa] considered the Dirichlet series
$\mathcal{D}_{F,G}(s)=\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$
[…]
@@
where $Z_{F}(s)$ is the spinor $L-$function attached to $F$. This gives a way of proving the meromorphic continuation of $Z_{F}(s)$ to $\mathbb{C}$.
[…]
@@
In this paper, we consider a Dirichlet series of the same type, for two cups forms $F,G$ of orthogonal groups of real signature $(2,n+2)$, which are eigenforms for their corresponding Hecke algebras. Here, we take $G$ to be a Maass lift of a Poincaré series. The Maass space in this setting has been defined by Gritsenko in [gritsenko_jacobi], Sugano in [sugano_lifting] and Krieg in [krieg_jacobi]. Our main goal is to show how the method of Kohnen and Skoruppa in [kohnen_skoruppa] can be extended in this setting and investigate the connection of this Dirichlet series with the standard $L-$function attached to $F$. The question of its analytic continuation and functional equation is quite different in nature and will be the subject matter of an upcoming paper.
@@
There is a variety of reasons why we should expect such a connection. First of all, the orthogonal groups of signature $(2,n+2)$ are isomorphic to classical groups for small $n$. For example, if $n=1$, there is an isomorphism with the symplectic and paramodular groups of degree 22 22. For $n=2$, there is an isomorphism with the Hermitian groups of degree 22 22, for $n=4$ with quaternionic groups of degree 22 22 and for $n=6$ with the linear groups of Cayley numbers of degree 22 22 (see for example [eisenstein_thesis, page 5]). Therefore, already the results of Kohnen and Skoruppa in [kohnen_skoruppa] and Gritsenko in [gritsenko] should indicate a connection in the $n=1,2$ case. Moreover, it should be noted that Gritsenko in [gritsenko_dirichlet_orthogonal] initiated the study of a Dirichlet series involving the Fourier coefficients of an orthogonal modular form and its connection with the standard $L-$function attached to it, following factorisation methods similar to [gritsenko]. Sugano in [sugano] extended Gritsenko’s result, proving an Euler product relation for a Dirichlet series involving twists of the Fourier coefficients by functions on an orthogonal group of lower rank. His work can be seen as an extension of Andrianov’s work in [andrianov] in the orthogonal setting. As we mentioned, our methods in this paper can be considered as an extension of the methods employed by Kohnen and Skoruppa in [kohnen_skoruppa] and therefore the result in [sugano] by Sugano is pivotal. Finally, results proved by Shimura in [shimura2004arithmetic] on the sets of solutions $\phi[x]=q$, where $\phi$ is a bilinear form and $q\in\mathbb{Q}^{\times}$, turn out to be crucial.
@@
$S_{0}=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},S_{1}=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
@@
$G^{*}_{K}=\{g\in SL_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},G_{K}=\{g\in SL_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
It turns out that $G_{\mathbb{R}}^{0}$, the connected component of the identity, acts on a suitable upper half plane, which we will call $\mathcal{H}_{S}\subset\mathbb{C}^{n+2}$, and a suitable factor of automorphy $j(\gamma,Z)$ can be defined. Let us now denote by $\Gamma_{S}$ the intersection of $G_{\mathbb{R}}^{0}$ with the stabiliser of the lattice $L_{1}=\mathbb{Z}^{n+4}$ and let $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ be an orthogonal cusp form with respect to $\Gamma_{S}$ (see Definitions 2.2 and 2.5).
If now $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$, we can write
@@
[…]
@@
[…]
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}V_{N}P_{k,D,r}(\tau,z)e(N\tau^{\prime}),$
[…]
$H(\xi)_{\mathbb{Q}}=\{g\in G_{\mathbb{Q}}^{*}\mid g\xi=\xi\}.$
Consider then the case when the class number $h:=\#(H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ equals 11 11, where $C=\{x\in G_{\mathbb{A}}^{*}\mid xL_{0}=L_{0}\}$. Here, the subscript A denotes the adelization of an algebraic linear group (see Section 6). Then, our main theorem 7.2 reads:
@@
With the notation and assumptions as above, let $\mathcal{P}$ denote a sufficiently large finite set of prime numbers. Define then the Dirichlet series
@@
We then have that $L_{\mathcal{P}}\left(\mathbf{1};s-k+\frac{n+3}{2}\right)\mathcal{D}_{F,P_{k,D,r},\mathcal{P}}(s)$ can be written as:
$A(\xi)L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\zeta_{\xi,\mathcal{P}}(s-k+n+1)\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases}.$
@@
Note here that the Theorem 7.2 covers the case $h>1$ as well. The rest of the paper (Section 8) is devoted in obtaining explicit examples where $h=1$. In those cases, we obtain a clear-cut Euler product connection between the specific Dirichlet series $\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}$ and $L_{\mathcal{P}}(F,s)$ (here we make a particular choice for $D$, i.e. $D=-q$).
We find all instances when this happens in the rank 11 11 case and we give the result in Theorem 8.1. Note that Theorem 8.1 recovers the case of Kohnen and Skoruppa in [kohnen_skoruppa] (we take $S=2$). What is more, it already produces new results related to $L-$functions attached to paramodular forms, as it is not clear how can one use the methods in [kohnen_skoruppa] to produce such results for paramodular forms directly. Finally, we provide at least one explicit example for each of the cases $n=2,4,6,8$, giving us Theorem 8.2. These examples can be related to other classical modular forms, such as Hermitian ($n=2$) and Quaternionic ($n=4$).
[…]
@@
Let $\mathbb{Z},\mathbb{Q},\mathbb{R},\mathbb{C}$ denote the usual rings of integers, rationals, real and complex numbers, respectively. For a complex number $z$, we denote by $e(z):=e^{2\pi iz}$. We will use the bracket notation $A[B]:=\overline{B}^{t}AB$ for suitably sized complex matrices $A,B$.
@@ 2. Preliminaries
Let $V$ denote a finite dimensional vector space over $\mathbb{Q}$. We start with the following definition:
@@
A $\mathbb{Z}-$lattice $\Lambda$ is a free, finitely generated $\mathbb{Z}-$module, which spans $V$ over $\mathbb{Q}$, together with a symmetric non-degenerate bilinear form $V\times V\longrightarrow\mathbb{Q}$.
In the following, let $V:=\mathbb{Q}^{n}$ and $L:=\mathbb{Z}^{n}$ with $n\geq 1$. Assume $S$ is an even integral positive definite symmetric matrix of rank $n\geq 1$. Even here means $S[x]\in 2\mathbb{Z}$ for all $x\in L$.
Let then $V_{0}:=\mathbb{Q}^{n+2}$ and consider
$S_{0}=\begin{pmatrix}0&0&1\\ 0&-S&0\\ 1&0&0\end{pmatrix}$
of real signature $(1,n+1)$. Define also the bilinear form
$\displaystyle\phi_{0}:V_{0}\times V_{0}$ $\displaystyle\longmapsto\mathbb{Q}$
$\displaystyle(x,y)$ $\displaystyle\longmapsto\frac{1}{2}x^{t}S_{0}y.$
Consider now a Witt decomposition of the form
$V_{0}=e_{n+1}\perp V\perp e_{1},$
where $e_{i}$ is the $i$’th standard basis vector. We then have that $\phi:=\phi\mid_{V\times V}$ is just $(x,y)\longmapsto-\frac{1}{2}x^{t}Sy$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}-$lattice with respect to $\phi$. Here, maximal means that the lattice $L$ is maximal among all $\mathbb{Z}-$lattices in $V$ on which $\phi[x]$ takes values in $\mathbb{Z}$, where we use the notation $\phi[x]:=\phi(x,x)$ for $x\in V$.
From [shimura2004arithmetic, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}-$maximal lattice in $V_{0}=\mathbb{Q}^{n+2}$.
Moreover, define $S_{1}=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$ and define the corresponding special orthogonal groups:
$G^{*}_{\mathbb{Q}}:=\{g\in\textup{SL}_{n+2}(\mathbb{Q})\mid g^{t}S_{0}g=S_{0}\},$
$G_{\mathbb{Q}}:=\{g\in\textup{SL}_{n+4}(\mathbb{Q})\mid g^{t}S_{1}g=S_{1}\}.$
We view $G^{*}_{\mathbb{Q}}$ as a subgroup of $G_{\mathbb{Q}}$ via the embedding $g\longmapsto\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}$.
We also use the notation $SO^{\phi_{0}}(V_{0})$ and in general $SO^{\phi}(V)$ for any quadratic space $(V,\phi)$.
@@
In particular, if we denote by $\mathcal{P}_{S}:=\{y=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\phi_{0}[y]>0\}$, we choose:
@@
$Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$
@@
$j(\gamma,Z):=-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta.$
@@
Let also $\Gamma^{+}(L_{0})=\Gamma(L_{0})\cap G^{*,0}_{\mathbb{R}}$. If then
@@
we have the following definition:
@@
Let $k\in\mathbb{Z}$. Then a holomorphic function $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ is called a modular form of weight $k$ with respect to the group $\Gamma_{S}$ if it satisfies the equation
@@
for all $\gamma\in\Gamma_{S}$ and $Z\in\mathcal{H}_{S}$. We will denote the set of such forms by $M_{k}(\Gamma_{S})$.
@@
Now, $F\in M_{k}(\Gamma_{S})$ admits a Fourier expansion of the form (cf. [eisenstein_thesis, page 25]):
$F(Z)=\sum_{r\in L_{0}^{*}}A(r)e(r^{t}S_{0}Z).$
where $Z\in\mathcal{H}_{S}$. It is then Koecher’s principle that gives us that $A(r)=0$ unless $r\in L_{0}^{*}\cap\overline{\mathcal{P_{S}}}$ (cf. [eisenstein_thesis, Theorem 1.5.2]). By [eisenstein_thesis, Theorem 1.6.23], we have the following definition for cusp forms.
@@
$F\in M_{k}(\Gamma_{S})$ is called a cusp form if it admits a Fourier expansion of the form
$F(Z)=\sum_{r\in L_{0}^{*}\cap\mathcal{P_{S}}}A(r)e(r^{t}S_{0}Z).$
We denote the space of cusp forms by $S_{k}(\Gamma_{S})$.
Finally, we give the definition of the Maass space, as this is defined in [krieg_jacobi, Section 5]:
@@
The Maass space $M_{k}^{*}(\Gamma_{S})$ consists of all $F\in M_{k}(\Gamma_{S})$, so that if their Fourier expansion is
@@
where $r=S_{0}^{-1}\rho$, with $\rho=\begin{pmatrix}m\\ \lambda\\ l\end{pmatrix}$.
[…]

### v2

# A Fourier–Jacobi Dirichlet series for cusp forms on orthogonal groups

@@ Abstract.
We investigate a Dirichlet series involving the Fourier–Jacobi coefficients of two cusp forms $F,G$ for orthogonal groups of signature $(2,n+2)$. In the case when $F$ is a Hecke eigenform and $G$ is a Maass lift of a Poincaré series, we establish a connection with the standard $L$-function attached to $F$. What is more, we find explicit choices of orthogonal groups, for which we obtain a clear-cut Euler product expression for this Dirichlet series. Through our considerations, we recover a classical result for Siegel modular forms, first introduced by Kohnen and Skoruppa, but also provide a range of new examples, which can be related to other kinds of modular forms, such as paramodular, Hermitian, and quaternionic.
@@ Acknowledgements.
The author would like to thank Prof. Thanasis Bouganis for suggesting this problem as well as for his continuous guidance and support. The author would also like to thank the anonymous referees for their useful comments and suggestions.
@@ 1. Introduction
Kohnen and Skoruppa in [13] considered the Dirichlet series
$\mathcal{D}_{F,G}(s)=\zeta(2s-2k+4)\sum_{N=1}^{\infty}\langle\phi_{N},\psi_{N}\rangle N^{-s},$
where $\phi_{N},\psi_{N}$ are the $N$th Fourier-Jacobi coefficients of two genus two Siegel cusp forms $F,G$ of integral weight $k\geq 0$ and $\langle\;,\;\rangle$ denotes the Petersson inner product on the space of Jacobi forms of index $N$ and weight $k$. They showed that this Dirichlet series has an analytic continuation to $\mathbb{C}$ and satisfies a functional equation. Moreover, in the case when $F$ is a Hecke eigenform and $G$ is in the Maass space, they deduced that
@@
where $Z_{F}(s)$ is the spinor $L$-function attached to $F$. In the case when $\phi_{1}$ does not vanish identically, this gives a way of proving the meromorphic continuation of $Z_{F}(s)$ to $\mathbb{C}$. Recently, Manickam in [16] showed that the first Fourier-Jacobi coefficient of a degree $2$ cuspidal Siegel eigenform does not vanish identically, hence this is actually not a restriction.
It was not too long after the work of Kohnen and Skoruppa that Gritsenko in [8] considered a same type Dirichlet series, however this time involving the Fourier–Jacobi coefficients of two Hermitian cusp forms of degree $2$. Gritsenko managed to prove that $\mathcal{D}_{F,G}(s)$ has an analytic continuation to $\mathbb{C}$ and that if $F$ is Hecke eigenform and $G$ is in the corresponding Maass space, $\mathcal{D}_{F,G}(s)$ is related to a degree $6$ $L$-function attached to $F$ ([8, (5.3)]). This $L$-function was first introduced by Gritsenko in [6], in his study of a Dirichlet series involving the Fourier coefficients of a Hermitian cuspidal eigenform of degree $2$. It is related to the exterior square representation of $\textup{GL}_{4}$ (see [19, Section 3.2.2] for example).
Even though the results in [13] and [8] are very similar in nature, the methods used by the authors are quite different. In particular, in [13], the authors prove their main result using a fundamental identity proved by Andrianov in [1]. This gives an Euler product expression for a certain Dirichlet series involving the Fourier coefficients of a Siegel cuspidal eigenform of degree $2$. In [8] on the other hand, Gritsenko manages to factorise the polynomials defining the $p$-factors of the degree $6$ $L$-function attached to a Hermitian cuspidal eigenform of degree $2$. This factorisation happens in some parabolic Hecke rings which contain elements acting on the Fourier–Jacobi coefficients.
@@
In this paper, we consider a Dirichlet series of the same type, attached to two cusp forms $F$ and $G$ for orthogonal groups of real signature $(2,n+2),\ n\geq 1$. We take $F$ to be a Hecke eigenform for the corresponding Hecke algebra and $G$ to be a Maass lift of a Poincaré series. The Maass space in this setting has been defined by Gritsenko in [7], Sugano in [28] and Krieg in [14]. Our main goal is to show how the method of Kohnen and Skoruppa in [13] can be extended in this setting and investigate the connection of this Dirichlet series with the standard $L$-function attached to $F$. The question of its analytic continuation and functional equation is quite different in nature and is the subject matter of our paper [20].
The main reason we expect such a connection is the accidental isogenies of orthogonal groups of signature $(2,n+2)$ with classical groups for small $n$. For example, for $n=1$, there is an isogeny with symplectic groups of degree $2$, for $n=2$, an isogeny with unitary groups of degree $2$, for $n=4$ with quaternionic groups of degree $2$ and for $n=6$ with the linear groups of Cayley numbers of degree $2$ (see for example [21, page 5]). Therefore, already the results of Kohnen and Skoruppa in [13] and Gritsenko in [8] indicate a connection in the $n=1,2$ case. We note here that the isogeny between $\textup{SO}(2,3)$ and $\textup{Sp}_{2}$ in particular was recently exploited by Pollack and Shah in [18] in order to extend the integral of Kohnen and Skoruppa to apply for arbitrary cuspidal automorphic representations of $\textup{PGSp}_{2}$.
@@
It should be noted that Gritsenko in [5] initiated the study of a Dirichlet series involving the Fourier coefficients of an orthogonal modular form and its connection with the standard $L$-function attached to it, following factorisation methods similar to [8]. Sugano in [27] extended Gritsenko’s result by proving an Euler product relation for a Dirichlet series involving twists of the Fourier coefficients by modular forms on a definite orthogonal group of lower rank. His work can be seen as an extension of Andrianov’s work in [1] in the orthogonal setting. As we mentioned, our methods in this paper can be considered as an extension of the methods employed by Kohnen and Skoruppa in [13], and therefore the result in [27] by Sugano is pivotal. Finally, results proved by Shimura in [24] on the sets of solutions $\phi(x,x)=q$, where $\phi$ is a bilinear form and $q\in\mathbb{Q}^{\times}$, turn out to be crucial.
We remark that modified Dirichlet series involving Fourier-Jacobi coefficients can also be related to Euler products. This happens in [12] for example, where the case of $\textup{GSp}_{2}\times\textup{GL}_{2}$ is treated. Jointly with Bouganis in [2], we partially extended this idea to $\textup{GU}(2,2)\times\textup{GL}_{2}$. The motivation is that these Dirichlet series admit integral representations of Rankin-Selberg type and therefore provide a way of proving the analytic properties of $L$-functions but also results on the algebraicity of their special values (see for example [3], where the result in [12] finds a profound application). The isogenies we mentioned above mean that we can consider these as integral representations for the twisted standard $L$-functions of $\textup{SO}(2,3)$ and $\textup{SO}(2,4)$ respectively.
@@
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\textup{ }S_{1}:=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
@@
$G^{*}_{K}=\{g\in\hbox{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},\textup{ }G_{K}=\{g\in\hbox{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
The connected component of the identity $G_{\mathbb{R}}^{0}$ acts on a suitable upper half plane, which we will call $\mathcal{H}_{S}\subset\mathbb{C}^{n+2}$, and a suitable factor of automorphy $j(\gamma,Z)$ can be defined. Let us now denote by $\Gamma_{S}$ the intersection of $G_{\mathbb{R}}^{0}$ with the stabiliser of the lattice $L_{1}=\mathbb{Z}^{n+4}$ and let $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ be an orthogonal cusp form of integral even weight $k>n/2+2$ with respect to $\Gamma_{S}$ (see Definitions 2.2 and 2.5). If now $Z=(\tau^{\prime},z,\tau)\in\mathcal{H}_{S}$, with $\tau,\tau^{\prime}\in\mathbb{C}$ and $z\in\mathbb{C}^{n+2}$, we can write
@@
and we call the functions $\phi_{m}(\tau,z)$ the Fourier–Jacobi coefficients of $F$. Now, if $G$ is another orthogonal cusp form with respect to some subgroup $\Gamma$, with $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$ ($\widetilde{\Gamma}_{S}$ here denotes the discriminant kernel), and with corresponding Fourier–Jacobi coefficients $\{\psi_{m}\}_{m=1}^{\infty}$, the object of interest is
@@
where $\langle\;,\;\rangle$ is a suitable inner product defined on the space of Fourier–Jacobi forms of certain weight and (lattice) index (see Definition 3.5). This series converges absolutely for $\textup{Re}(s)>k+1$ and represents a holomorphic function in this domain (see Lemma 6.1). Our aim is to find an Euler product expression for this Dirichlet series.
We will now present our main Theorem 8.2 in a simplified form. Let $F$ be an eigenform for some appropriately defined Hecke algebras (see Theorem 8.2) and with Fourier–Jacobi expansion as above. Let $\sigma$ denote the bilinear form induced by $S$, $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$ (see Definition 5.1) and define
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P_{k,D,r})(\tau,z)e(N\tau^{\prime}),$
where $P_{k,D,r}$ is a specific Fourier–Jacobi form which reproduces the Fourier coefficients of any Fourier–Jacobi form, under the Petersson inner product (see Section 5). Also, $V_{N}$ denotes the index raising operator of Section 4, acting on Fourier–Jacobi forms. Let $L_{0}=\mathbb{Z}^{n+2}$ and $L_{0}^{*}=S_{0}^{-1}L_{0}$. Assume now it exists $\xi\in V_{0}:=\mathbb{Q}^{n+2}$ such that $\{\xi^{t}S_{0}x\mid x\in L_{0}\}=\mathbb{Z}$ and $S_{0}[\xi]=-2D/q$, where $q$ denotes the level of $S$ (see Definition 2.4) and further assume $A(\xi)\neq 0$ for the Fourier coefficient of $F$. Consider then the algebraic subgroup of $G_{\mathbb{Q}}^{*}$, defined by
$H(\xi)_{\mathbb{Q}}=\left\{g\in G_{\mathbb{Q}}^{*}\mid g\xi=\xi\right\}.$
We consider the case when the class number $h:=\#(H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ equals $1$, where $C=\{x\in G_{\mathbb{A}}^{*}\mid xL_{0}=L_{0}\}$. Here, the subscript A denotes the adelization of an algebraic linear group (see Section 7). Then, our main Theorem 8.2 reads:
@@
With the notation and assumptions as above, let $\mathcal{P}$ denote a sufficiently large (specified) finite set of prime numbers. Define then the Dirichlet series
@@
We then have that $L_{\mathcal{P}}\left(\mathbf{1};s-k+(n+3)/2\right)\mathcal{D}_{F,P_{k,D,r},\mathcal{P}}(s)$ can be written as:
$A(\xi)L_{\mathcal{P}}\left(F;s-k+(n+2)/2\right)\zeta_{\xi,\mathcal{P}}(s-k+n+1)\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases}.$
@@
Note that in Theorem 8.2, we cover the case $h>1$ as well. The rest of the paper (Section 9) is devoted in obtaining explicit examples where $h=1$. In those cases, we obtain a clear-cut Euler product expression for the specific Dirichlet series $\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}(s)$ (here we make a particular choice for $D$, i.e. $D=-q$ and take $\xi=(1,\textbf{0},1)^{t}$).
We find all instances when this happens in the rank $1$ case and we give the result in Theorem 9.2. Note that in Theorem 9.2, we recover the case of Kohnen and Skoruppa in [13] for our choice of $D$ (we take $S=2$), up to a finite number of Euler factors. What is more, we already produce new results related to $L$-functions attached to paramodular forms, as it is not clear how can one use the methods in [13] to produce such results for paramodular forms directly. Finally, in Theorem 9.3, we provide at least one explicit example for each of the cases $n=2,4,6,8$. These examples can be related to modular forms for other classical groups, such as Hermitian ($n=2$) and quaternionic ($n=4$).
The structure of the paper is as follows: In Sections 2 and 3 we give the definitions of the main objects of consideration, i.e. orthogonal groups, modular forms and Fourier–Jacobi forms. In Section 4, we calculate the action of the adjoint of the level-raising operator $V_{N}$ with respect to the Petersson inner product of Fourier–Jacobi forms. In Section 5, we define Poincaré series and in Section 6, we define the Dirichlet series of interest and prove some preliminary results. Section 7 is devoted in showing how results of Shimura in [24] and [26] can be used in our setting and in Section 8, we use Sugano’s theorem in [27] to deduce our main Theorem 8.2. Finally, in Section 9, we obtain clear-cut Euler product expressions for some specific orthogonal groups.
@@
We denote the space of $m\times n$ matrices with coefficients in a ring $R$ with $\textup{M}_{m,n}(R)$. If $n=m$, we often use the notation $\textup{M}_{n}(R)$. By $1_{n}$, we denote the $n\times n$ identity matrix. For any matrix $M\in\textup{M}_{n}(R)$, we denote by $\det(M),\textup{ tr}(M)$ the determinant and trace of $M$ respectively. By $\textup{GL}_{n}(R)$, we denote the matrices in $M_{n}(R)$ with non-zero determinant and by $\textup{SL}_{n}(R)$ the matrices with determinant $1$. For any vector $v$, we denote by $v^{t}$ its transpose. We also use the bracket notation $A[B]:=\overline{B}^{t}AB$ for suitably sized complex matrices $A,B$. By $\textup{diag}(A_{1},A_{2},\cdots,A_{n})$, we will denote the block diagonal matrix with the matrices $A_{1},A_{2},\cdots,A_{n}$ in the diagonal blocks. For a complex number $z$, we denote by $e(z):=e^{2\pi iz}$. Finally, let $\zeta(s)$ denote the Riemann zeta function and $\Gamma(s):=\int_{0}^{\infty}t^{s-1}e^{-t}\hbox{d}t$ denote the Gamma function.
@@ 2. Preliminaries
Let $V$ denote a finite dimensional vector space over $\mathbb{Q}$. We start with the following definition.
@@
A $\mathbb{Z}$-lattice $\Lambda$ is a free, finitely generated $\mathbb{Z}$-module, which spans $V$ over $\mathbb{Q}$.
In the following, let $V:=\mathbb{Q}^{n}$ and $L:=\mathbb{Z}^{n}$ with $n\geq 1$. Assume $S$ is an even integral positive definite symmetric matrix of rank $n$. Here, even means $S[x]\in 2\mathbb{Z}$ for all $x\in L$. We define
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\ S_{1}:=\begin{pmatrix}&&1\\ &S_{0}&\\ 1&&\end{pmatrix}$
of real signatures $(1,n+1)$ and $(2,n+2)$ respectively. Let also $V_{0}:=\mathbb{Q}^{n+2}$ and $V_{1}:=\mathbb{Q}^{n+4}$ and consider the quadratic spaces $(V_{0},\phi_{0}),\ (V_{1},\phi_{1})$, where
$\displaystyle\phi_{i}:V_{i}\times V_{i}$ $\displaystyle\longmapsto\mathbb{Q}$
$\displaystyle(x,y)$ $\displaystyle\longmapsto\frac{1}{2}x^{t}S_{i}y,$
for $i=0,1$. We then have that $\phi:=\phi_{0}\mid_{V\times V}$ is just $(x,y)\longmapsto-x^{t}Sy/2$, and we make the assumption that $L=\mathbb{Z}^{n}$ is a maximal $\mathbb{Z}$-lattice with respect to $\phi$. Here, maximal means that the lattice $L$ is maximal among all $\mathbb{Z}$-lattices in $V$ on which $\phi[x]$ takes values in $\mathbb{Z}$, where we use the notation $\phi[x]:=\phi(x,x)$ for $x\in V$.
From [24, Lemma 6.3], we then obtain that $L_{0}:=\mathbb{Z}^{n+2}$ is a $\mathbb{Z}$-maximal lattice in $V_{0}$.
If now $K\supset\mathbb{Q}$ is a field, we define the corresponding special orthogonal groups of $K$-rational points via
$G^{*}_{K}:=\{g\in\textup{SL}_{n+2}(K)\mid g^{t}S_{0}g=S_{0}\},$
$G_{K}:=\{g\in\textup{SL}_{n+4}(K)\mid g^{t}S_{1}g=S_{1}\}.$
We view $G^{*}_{K}$ as a subgroup of $G_{K}$ via the embedding $g\longmapsto\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}$.
@@
In particular, if we denote by $\mathcal{P}_{S}:=\{y^{\prime}=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\phi_{0}[y^{\prime}]>0\}$, we choose
@@
$Z\longmapsto g\langle Z\rangle=\frac{-\frac{1}{2}S_{0}[Z]b+AZ+c}{-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta}$ (1)
@@
$j(g,Z):=-\frac{1}{2}S_{0}[Z]\gamma+d^{t}Z+\delta.$
@@
Let also $\Gamma^{+}(L_{0})=\Gamma(L_{0})\cap G^{*,0}_{\mathbb{R}}$. Let now
@@
and
$\widetilde{\Gamma}_{S}:=\{M\in\Gamma_{S}\mid M\in 1_{n+4}+\textup{M}_{n+4}(\mathbb{Z})S_{1}\},$
the discriminant kernel.
@@
Let $k\in\mathbb{Z}$ and $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$ a subgroup of finite index. A holomorphic function $F:\mathcal{H}_{S}\longrightarrow\mathbb{C}$ is called a modular form of weight $k$ with respect to $\Gamma$ if it satisfies the equation
@@
for all $\gamma\in\Gamma$ and $Z\in\mathcal{H}_{S}$. We will denote the set of such forms by $M_{k}(\Gamma)$.
@@
Now, if $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$, $F\in M_{k}(\Gamma)$ admits a Fourier expansion of the form (see [28, (5.10)])
$F(Z)=\sum_{r\in L_{0}^{*}}A(r)e(r^{t}S_{0}Z),$ (2)
where $Z\in\mathcal{H}_{S}$. It is then Koecher’s principle that gives us that $A(r)=0$ unless $r\in L_{0}^{*}\cap\overline{\mathcal{P_{S}}}$ ($\overline{\mathcal{P}_{S}}$ denotes the closure of $\mathcal{P_{S}}$, see [21, Theorem 1.5.2]). By [21, Theorem 1.6.23], and because $L$ is maximal, we have the following definition for cusp forms.
@@
If $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$, $F\in M_{k}(\Gamma)$ is called a cusp form if it admits a Fourier expansion of the form
$F(Z)=\sum_{r\in L_{0}^{*}\cap\mathcal{P_{S}}}A(r)e(r^{t}S_{0}Z),$
We denote the space of cusp forms by $S_{k}(\Gamma)$.
Finally, we give the definition of the Maass space, due to Krieg in [14, Section 5].
@@
Let $\widetilde{\Gamma}_{S}\leq\Gamma\leq\Gamma_{S}$. The Maass space $M_{k}^{*}(\Gamma)$ consists of all $F\in M_{k}(\Gamma)$, so that if their Fourier expansion is
@@
where $r=S_{0}^{-1}\rho$, with $\rho=\begin{pmatrix}m&\lambda&l\end{pmatrix}^{t}$. We also write $S_{k}^{*}(\Gamma)$ for the Maass space consisting of cusp forms.
3. Fourier–Jacobi Forms
In this section, we will define Fourier–Jacobi forms of lattice index. We mainly follow Mocanu’s thesis in [17] and Krieg in [14].